• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 16:00:51 on August 18, 2006

    Entry number 180709

    keyword=Day shift summary

    8:00 Shift started with poor FPP efficiency and missing GUI information on Hall A tools and LeCroy HV screens. Suspected gas issues on FPP and tried to reboot IOCHA22 for tools. Reboot seemed OK but information still did not show. Rebooted multiple times, could also not see BPM. Finally closed tools and hlamain, re-opened them and infomation showed up.

    In parallel, Mark Jones took an access to inspect chambers in Hall while Alexander, Rob and techs checked gas shed. After changing "stuff" outside, we are now using an argon bottle instead of dewar adn the argon pressure is now 2400 (instead of 80). Nothing wrong in Hall noticed by Mark Jones.

    9:00 waiting for beam (MCC doing Mott measurements) to see if efficiency is now OK.

    10:00 still waiting for beam. MCC did Mott measurements (see ELOG) and got 83% beam polarization. Since then the MCC has been trying to restore beam.

    10:16 First trickles of beam in hall.

    10:30 FPP efficiency is still low. [See run 2870.] If the low efficiency is due to bad chamber gas, it may take several hours to flush sufficiently.

    10:40 RC decides to do a harp scan. Lower current to 5 uA. Raster is still on. First attempts yield only a suggestion that there is a signal. Move to carbon target and take out radiator. Switch off raster.

    11:35 put in BeO. Suspect that the normal carbon target is actually empty. Spot gives size of about 2 mm by 1.5 mm FWHM. MCC claims sigma to be 1.3 mm and 1.8 mm.

    12:50 Jay Benesch has improved the spot to .5 mm by 1mm FWHM on the harp. Spot shows FWHM to be about .7mm by .6 mm.

    13:30 Put LD2 15 cm target in and radiator 5% back. Asked for 10 uA beam with raster on. Still waiting for improvement in FPP efficiency.

    2:00 Increased FPP HV to 1825, per Mark Jones request for first set of FPP HV (left second set at 1725). Run 2876.

    2:10 Front FPP chambers seemed to do a little better. Increased all chambers to 1825 V. Efficiency improved somwewhat (run 2877).

    2:40 Setting up for Bullseye scan. Going to carbon optics target, taking radiator out.

    CODA (L/R) Harp position

    2878 21222 2258 0,0 2879 21223 2259,2260,2261 -2,0 2880 21224 2262 +2,0 2881 21225 2263 +2,+2 2882 21226 2265 0,+2 2883 21227 2266 0,-2 2884 21228 2267 -2,-2 2885 21229 2268 0,0